### v2

# Exploring Multiple Strategies to Improve Multilingual Coreference Resolution in CorefUD

Abstract
Coreference resolution, the task of identifying expressions in text that refer to the same entity, is a critical component in various natural language processing (NLP) applications. This paper presents our end-to-end neural coreference resolution system, utilizing the CorefUD 1.1 dataset, which spans 17 datasets across 12 languages. Our model is based on the end-to-end neural coreference resolution system by [1]. We first establish strong baseline models, including monolingual and cross-lingual variations, and then propose several extensions to enhance performance across diverse linguistic contexts. These extensions include cross-lingual training, incorporation of syntactic information, a Span2Head model for optimized headword prediction, and advanced singleton modeling. We also experiment with headword span representation and long-document modeling through overlapping segments. The proposed extensions, particularly the heads-only approach, singleton modeling, and long document prediction, significantly improve performance across most datasets. We also perform zero-shot cross-lingual experiments, highlighting the potential and limitations of cross-lingual transfer in coreference resolution. Our findings contribute to the development of robust and scalable coreference systems for multilingual coreference resolution. Finally, we evaluate our model on the CorefUD 1.1 test set and surpass the best model from the CRAC 2023 shared task of comparable size by a large margin. Our model is available on GitHub: https://github.com/ondfa/coref-multiling
Keywords coreference resolution  $\cdot$ cross-lingual model  $\cdot$ Transformers  $\cdot$ end-to-end model
1 Introduction
Coreference resolution is the task of identifying language expressions that refer to the same real-world entity (antecedent) within a text. These coreferential expressions can sometimes appear within a single sentence, but often, they are spread across multiple sentences. In some challenging cases, it is necessary to consider the entire document to determine whether two expressions refer to the same entity. The task can be divided into two main subtasks: identifying entity mentions and grouping these mentions based on the real-world entities they refer to. Coreference resolution is closely related to anaphora resolution, as discussed in [2]
Historically, coreference resolution was a standard preprocessing step in various natural language processing (NLP) tasks, such as machine translation, summarization, and information extraction. Although recent large language models have achieved state-of-the-art results in coreference resolution, they are expensive to train and deploy, and traditional (discriminative) approaches remain competitive. Expressing this task in natural language is challenging, and to the best of our knowledge, there have been no successful attempts to utilize large chatbots (like ChatGPT-4) to achieve superior results.
Coreference resolution becomes particularly challenging in low-resource languages. One strategy to address this challenge is to train a multilingual model on datasets from multiple languages, thereby transferring knowledge from resource-rich languages to those with fewer resources. However, a significant challenge with this approach lies in the differences in annotations across available corpora. The CorefUD initiative [3] tries to harmonize the datasets and create one annotation scheme for coreference in multiple languages similarly to Universal Dependencies [4] for syntactic annotations.
This paper describes our approach to multilingual coreference resolution. The task is based on the CorefUD dataset [3]. The CorefUD 1.1 corpus comprises 17 different datasets spanning 12 languages, all within a harmonized annotation scheme. As CorefUD is intended to extend Universal Dependencies with coreference annotations, all datasets within CorefUD are treebanks. For some languages, dependency annotations were provided by human annotators, while for others, these annotations were generated automatically using a parser. Coreference annotations are built upon these dependencies, meaning that mentions are represented as subtrees in the dependency tree and can be captured by their heads. In some datasets, there are non-treelet mentions—mentions that do not form a single subtree—but even for these, a single headword is selected. Notable differences exist between the datasets, with one of the most prominent being the presence of singletons. Singletons are clusters that contain only one mention and do not participate in any coreference relation, yet they are still annotated as mentions. For further details on the dataset, refer to [3] or [5]. We use the primary evaluation metric from CRAC 2023 Multilingual Coreference Resolution Shared Task. The metric ignores singletons. Building upon the baseline model proposed by [1], we introduce several novel extensions aimed at enhancing the performance of coreference resolution across multiple languages and datasets within the CorefUD collection. Our primary goal is to develop a universal model capable of handling the diverse and complex nature of these datasets effectively.
First, recognizing the challenge posed by small dataset sizes in the CorefUD collection, we propose a cross-lingual training approach. By pretraining our model on a concatenated dataset that includes all available training data across languages, we aim to improve the model’s ability to generalize across languages. This approach is particularly beneficial for low-resource languages, where training large models from scratch is impractical due to the limited data available.
Next, we incorporate syntactic information into our model to leverage the dependency structures inherent in the CorefUD datasets. This extension adds syntax to the token representations by encoding their paths to the ROOT in the dependency tree, thereby enriching the model’s understanding of syntactic relationships, which are crucial for accurate mention detection and coreference resolution.
In response to the evaluation metrics used in the CRAC 2022 Shared Task, we also propose the Span2head model, which simplifies span-based mention representation to headword prediction. This adjustment aligns the model’s output more closely with the evaluation criteria, where we consider two mentions the same if they share the same head.
Additionally, we explore the potential of head representations as a simplified approach to coreference resolution. By modeling mentions solely based on their syntactic heads, we reduce the computational complexity from quadratic to linear, making the model more efficient and less prone to errors, particularly in the case of long and complex mentions.
We further address the issue of singleton—mentions by introducing mechanisms to incorporate these into the training process. By modeling singletons explicitly, we ensure that valuable training data from singleton-rich datasets is not discarded, thus improving the model’s robustness and accuracy across different datasets.
Finally, to overcome the limitations of short sequence lengths in XLM-R, we propose an approach that utilizes overlapping segments with a cluster merging algorithm. This method ensures that coreference chains spanning multiple segments are correctly identified and merged, even when document segmentation is necessary due to memory constraints.
Through these extensions, our model aims to advance the state of multilingual coreference resolution by addressing the specific challenges posed by the diverse datasets in the CorefUD collection, ultimately contributing to more accurate and generalizable coreference systems.
2 Related Work
2.1 End-to-end Neural Coreference Resolution Models
The first end-to-end neural coreference resolution system was introduced by [6], and many subsequent neural coreference resolution systems are based on their model.
The model estimates the probability $P(y_{i}|D)$ of a mention $i$ corefering with the antecedent $y_{i}$ in a document $D$. Since the model adopts the end-to-end approach, the mentions are identified together with the coreference links. We consider every continuous sequence of words as a mention $i$. Therefore, we work with $N=\frac{T(T+1)}{2}$ possible mentions, where $T$ is the number of words in a document $D$.
We model the score of a mention $i$ corefering with an antecedent $y_{i}$ as a combination of two types of scores $s_{m}(i)$ and $s_{a}(i,y_{i})$. The $s_{m}$ is a score of a sequence of words (spans) $i$ being a mention. The $s_{a}(i,y_{i})$ score is the score of a span $y_{i}$ being an antecedent of span $i$. The scores are combined as a sum of $s_{m}(i)$, $s_{m}(y_{i})$ and $s_{a}(i,y_{i})$ as follows:
$s(i,y_{i})=\begin{cases}0&y_{i}=\epsilon\\ s_{m}(i)+s_{m}(y_{i})+s_{a}(i,y_{i})&y_{i}\neq\epsilon\\ \end{cases},$ (1)
where $\epsilon$ is an empty (also referred as dummy) antecedent. If a mention candidate is aligned to the dummy antecedent and it is not an antecedent of any other mention (so the mention should be in the cluster with dummy antecedent only), it is not considered a valid mention. As a result, every valid mention has to be in a cluster with at least one other mention, so the model cannot predict singletons. Both scores $s_{m}(i)$ and $s_{a}(i,y_{i})$ are estimated with a feed-forward neural network over the BERT-based encoder.
The probability of an antecedent $y_{i}$ can be expressed as the $softmax$ normalization over all possible antecedents $y^{\prime}\in Y(i)$ for a mention $i$:
$P(y_{i}|D)=\frac{\exp(s(i,y_{i})}{\sum_{y^{\prime}\in Y(i)}\exp(s(i,y^{\prime})}$ (2)
The formula for all antecedents uses a product of multinomials of all individual antecedents:
$P(y_{1},...,y_{N}|D)=\prod_{i=1}^{N}P(y_{i}|D)$ (3)
In the training phase, we maximize the marginal log-likelihood of all correct antecedents:
$J(D)=\log\prod_{i=1}^{N}\sum_{\hat{y}\in Y(i)\cap\texttt{GOLD}(i)}P(\hat{y})$ (4)
where GOLD($i$) is the set of spans (mentions) in the training data that are antecedents of mention $i$.
The schema of the model is shown in Figure 1.
(a) Mention Ranking
(b) Antecedent Ranking
Figure 1: Architecture of the model from [6]
This end-to-end model has proven to be highly successful, especially on datasets where singletons are not annotated (e.g., CoNLL 2012/OntoNotes), a scenario where traditional two-stage models tend to struggle. The model was later extended by [7] to the higher-order coreference resolution (CR) through iterative refining of span representations with a gated attention mechanism. They further optimize pro model speed by pruning the mentions with simple scorer at the first step and rescoring only top $k$ spans with the precise scorer and the second step. [8] uses the model from [7] but with BERT as an encoder.
While many higher-order CR approaches have been proposed, [1] demonstrated that their impact is marginal when a strong encoder is used, as seen in their experiments with SpanBERT [9].
2.2 Word-level Coreference Resolution
[10] proposed a word-level coreference resolution approach based on the model by [6]. Their method reduces the computational complexity by reducing the mention space. Instead of iterating over all possible spans they first map each gold mention to a single word—specifically, the headword, which they select based on the syntax tree (as the node closest to the root of the tree). Antecedent prediction is then performed on the word level. In the next step, they use a span extraction model to predict the original spans from headwords. This model is trained to classify each word in the sentence together with a headword to decide, whether the word is the start token or end token of the span corresponding to the headword. During training the span extraction model is trained simultaneously with with the antecedent prediction as another classification head.
Later [11] proposed a modification in headword selection, suggesting that when a coordinate conjunction appears within an entity, it should be made the head to resolve span ambiguity. For example, in the sentence "Tom and Mary are playing," the entities "Tom" and "Tom, and Mary" would typically share the same head in the original approach. Their modification eliminates this ambiguity.
2.3 CRAC Shared Task on Multilingual Coreference Resolution
CRAC Shared Task on Multilingual Coreference Resolution (CRAC-coref) [12] is an annual shared task that began in 2022 and is built upon the CorefUD collection. In fact, this paper is an extension of CRAC22-coref [13] participant system [14], which is based on [15]
Best-performing system in CRAC 2022 was submitted by [16]. It is a two-stage model with a joined BERT-like encoder. In the mention detection stage, they use extended BIO [^1] BIO stands for Beginning, Inside, and Outside. It is a tagging format to label each token of a text as being the beginning (B), inside (I), or outside (O) of an entity. tagging via stack manipulation instructions to solve overlapping mentions. Coreference resolution is then done by antecedent prediction head, a component responsible for identifying the most likely antecedent of each mention, which is very similar to the one used in end-to-end neural CR [6]. They use a concatenation of start and end tokens as mention representation and classify mention pairs to find the best antecedent of each mention. The schema of the CorPipe model is shown in Figure 2.
Figure 2: Schema of CorPipe model (from [16])
Other cross-lingual experiments include Portuguese by learning from Spanish [17]; Spanish and Chinese relying on an English corpus [18]; and Basque based on an English corpus as well [19]. All these approaches employ neural networks, and they transfer the model via cross-lingual word embeddings.
3 Datasets & Metrics
For our experiments, we use the CorefUD 1.1 dataset. The collection consists of 17 datasets in 12 different languages. Table 1 provides a summary of these datasets and their respective sizes.
3.1 Dataset Differences
As highlighted by [3] significant differences remain among the individual datasets within CorefUD, not only in data distribution but also in annotation practices. In this section, we discuss these differences that are particularly relevant to our model and experiments.
3.1.1 Empty Nodes
Some datasets contain empty nodes. These are virtual nodes added to a syntactic tree during the annotation process (and also into the text). These nodes refer to unexpressed entities; entities that are not explicitly mentioned in the text but are referenced. Adding empty nodes can make it hard for a general-purpose pre-trained model to to accurately interpret the meaning of the text. We cannot discard them from the datasets because most of them are part of a coreference cluster. Entity mentions that span only the empty nodes are called zero mentions.
3.1.2 Singletons
Another key difference between the datasets lies in the annotation of singletons. Singletons are entities with only a single mention throughout the document so these mentions have no coreference relations. In some datasets, singletons are annotated, while in others they are not. That means that when there is an entity with only a single mention, it is not considered an entity during annotation. Table 2 provides the percentage of singleton entities across the datasets (entities with a length of 1).
3.2 Dataset Lengths
Tables 1, 2, and 3 show basic statistics of all the datasets in CorefUD collection. All these tables are taken from the CorefUD paper [5]. Table 1 shows the basic sizes of the datasets and portion of the train/dev/test split. Column empty shows the number of empty nodes. Table 2 shows statistics about entities. The length of an entity means the number of its mentions thus entities with length 1 are singletons. Table 3 shows statistics of mentions. Here the length of a mention means the number of its words. The last three columns show the percentage of specific mentions. w/empty are mentions with at least one empty node, w/gap are mentions with at least one gap e.g. discontinuous mentions. non-tree are mentions which do not form a single subtree of the syntax tree.
Table 1: Dataset Statistics (taken from [3]). words is the number of non- empty UD nodes. empty is the number of empty UD nodes.
total size division [%]
CorefUD dataset docs sents words empty train dev test
Catalan-AnCora 1298 13,613 429,313 6,377 77.5 11.4 11.1
Czech-PCEDT 2312 49,208 1,155,755 35,844 80.9 14.2 4.9
Czech-PDT 3165 49,428 834,720 22,389 78.3 10.6 11.1
English-GUM 195 10,761 187,416 99 78.9 10.5 10.6
English-ParCorFull 19 543 10,798 0 81.2 10.7 8.1
French-Democrat 126 13,057 284,883 0 80.1 9.9 10.0
German-ParCorFull 19 543 10,602 0 81.6 10.4 8.1
German-PotsdamCC 176 2,238 33,222 0 80.3 10.2 9.5
Hungarian-KorKor 94 1,351 24,568 1,988 79.2 10.3 10.5
Hungarian-SzegedKoref 400 8,820 123,968 4,857 81.1 9.6 9.3
Lithuanian-LCC 100 1,714 37,014 0 81.3 9.1 9.6
Norwegian-BokmaalNARC 346 15,742 245,515 0 82.8 8.8 8.4
Norwegian-NynorskNARC 394 12,481 206,660 0 83.6 8.7 7.7
Polish-PCC 1828 35,874 538,885 470 80.1 10.0 9.9
Russian-RuCor 181 9,035 156,636 0 78.9 13.5 7.6
Spanish-AnCora 1356 14,159 458,418 8,112 80.0 10.0 10.0
Turkish-ITCC 24 4,733 55,341 0 81.5 8.8 9.7
Table 2: Entity Statistics (taken from [3]). The total number of entities and the average number of entities per 1000 tokens in the running text. The maximum and average entity “length”, i.e., number of mentions of the entity. Distribution of entity lengths.
Entities distribution of lengths
CorefUD dataset total per 1k length 1 2 3 4 5+
count words max avg. [%] [%] [%] [%] [%]
Catalan-AnCora 18,030 42 101 3.5 2.6 54.0 18.1 8.6 16.8
Czech-PCEDT 52,721 46 236 3.3 6.6 59.6 14.7 6.3 12.7
Czech-PDT 78,747 94 175 2.4 40.8 35.1 10.4 4.9 8.9
English-GUM 27,757 148 131 1.9 73.8 14.2 4.9 2.2 4.9
English-ParCorFull 202 19 38 4.2 6.9 54.0 13.9 5.9 19.3
French-Democrat 39,023 137 895 2.0 81.6 10.7 3.0 1.4 3.3
German-ParCorFull 259 24 43 3.5 6.2 64.9 11.6 5.0 12.4
German-PotsdamCC 3,752 113 15 1.4 76.5 13.9 5.0 1.8 2.7
Hungarian-KorKor 1,134 46 41 3.6 0.9 55.1 17.1 9.0 17.9
Hungarian-SzegedKoref 5,182 42 36 3.0 8.0 51.1 19.0 9.1 12.9
Lithuanian-LCC 1,224 33 23 3.7 11.2 45.3 11.8 8.2 23.5
Norwegian-BokmaalNARC 53,357 217 298 1.4 89.4 5.4 1.9 1.0 2.4
Norwegian-NynorskNARC 44,847 217 84 1.4 88.7 5.5 2.1 1.1 2.7
Polish-PCC 127,688 237 135 1.5 82.6 9.8 2.9 1.4 3.2
Russian-RuCor 3,636 23 141 4.5 3.3 53.7 15.6 6.9 20.5
Spanish-AnCora 20,115 44 110 3.5 3.3 53.7 17.0 8.8 17.1
Turkish-ITCC 690 12 66 5.3 0.0 39.0 19.6 10.7 30.7
Table 3: Non-singleton mention Statistics (taken from [3]). The total number of mentions and the average number of mentions per 1000 words of running text. The maximum and average mention length, i.e., number of non-empty nodes in the mention. Distribution of mention lengths.
mentions distribution of lengths mention type CorefUD dataset total per 1k length 0 1 2 3 4 5+ w/empty w/gap non-tree count words max avg. [%] [%] [%] [%] [%] [%] [%] [%] [%] Catalan-AnCora 62,417 145 141 4.8 10.2 28.2 21.7 7.9 5.3 26.8 12.4 0.0 3.7 Czech-PCEDT 168,138 145 79 3.6 19.5 29.9 16.8 8.6 4.1 20.9 25.7 0.8 8.4 Czech-PDT 154,983 186 99 3.1 10.7 39.6 20.3 9.1 4.2 16.1 13.7 1.3 2.2 English-GUM 32,323 172 95 2.6 0.0 56.5 19.5 8.0 3.8 12.1 0.0 0.0 1.2 English-ParCorFull 835 77 37 2.1 0.0 59.6 23.4 6.3 3.4 7.3 0.0 0.6 0.6 French-Democrat 46,487 163 71 1.7 0.0 64.2 21.8 6.3 2.4 5.3 0.0 0.0 2.0 German-ParCorFull 896 85 30 2.0 0.0 64.8 17.5 6.2 4.0 7.4 0.0 0.3 1.5 German-PotsdamCC 2,519 76 34 2.6 0.0 34.8 32.4 15.6 6.4 10.9 0.0 6.3 3.8 Hungarian-KorKor 4,103 167 42 2.2 30.8 20.6 23.2 9.8 4.8 10.8 35.3 0.6 5.5 Hungarian-SzegedKoref 15,165 122 36 1.6 15.1 37.4 32.5 10.2 2.6 2.2 15.2 0.4 1.1 Lithuanian-LCC 4,337 117 19 1.5 0.0 69.1 16.6 11.1 1.2 2.0 0.0 0.0 4.3 Norwegian-BokmaalNARC 26,611 108 51 1.9 0.0 74.5 9.7 6.1 2.1 7.6 0.0 0.6 1.5 Norwegian-NynorskNARC 21,847 106 57 2.1 0.0 70.2 10.1 7.7 2.8 9.2 0.0 0.4 1.4 Polish-PCC 82,804 154 108 2.1 0.3 68.7 14.9 5.2 2.7 8.2 0.5 1.0 4.7 Russian-RuCor 16,193 103 18 1.7 0.0 69.1 16.3 6.6 3.5 4.6 0.0 0.5 1.4 Spanish-AnCora 70,663 154 101 4.8 11.4 31.6 18.8 7.2 4.5 26.3 14.0 0.0 0.3 Turkish-ITCC 3,668 66 25 1.9 0.0 67.3 17.5 5.8 2.9 6.6 0.0 0.0 1.9
3.3 Evaluation Metrics
For evaluation, we employ the official metric for CRAC 2023 Multilingual Coreference Resolution shared task implemented via the CoreUD scorer [^2] https://github.com/ufal/corefud-scorer. The metric is a modification of the CoNLL score for coreference resolution, averaging three metrics: CEAF-e, $B^{3}$, and MUC. All these three metrics are described later in this section. The main modification of the CorefUD scorer lies in using the head matching. In head matching the gold and system mentions are considered identical if and only if they have the same syntactic head. Another important aspect is that the primary metric ignores singletons. The CorefUD scorer also provides additional metrics, such as BLANC [20] and LEA [21]. To assess the quality of mention matching while ignoring the assignment of mentions to coreferential entities, it uses the MOR score (mention overlap ratio) [13].
3.3.1 A Link-Based Metric: The MUC Score
The MUC official scorer [22] introduced a link-based metric. A link-based metric measures the extent to which the links in the response match the links in the key. For example, recall is computed by summing up the correctly recalled links for each coreference chain in the key and then dividing by the total number of correct links in the key. The number of missing links—the links found in the key entities but not in the response entities—is computed by counting the number of partitions of key K induced by response R, as follows:
$\text{Recall}_{\text{MUC}}=\frac{\sum_{i}|K_{i}|-|\mathcal{P}(K_{i},R)|}{\sum_{i}|K_{i}|-1}$ (5)
where $\mathcal{P}(K_{i},R)$ is the partition function, which returns all the partitions of key entity $K_{i}$ with respect to a system’s response $R$. Coreference chains are formed by linking each mention in a cluster to its antecedent, creating a sequence of references. Precision is computed by summing up the correct links in each coreference chain in the response and dividing that by the total number of links in the response. Essentially, precision evaluates how many of the predicted coreference links accurately match the gold standard. This is done by swapping key and response in the formula, where the system’s response is compared against the gold coreference chain, and the number of correct links is divided by the total predicted links.
3.3.2 A Mention-Based Metric: $B^{3}$
One problem with the MUC score is that, by definition, it only scores a system’s ability to identify links between mentions; its ability to recognize that a mention does not belong to any coreference chain—that is, its ability to classify a mention as a singleton—does not get any reward. The $B^{3}$ metric [23] was proposed to correct this problem. It does this by computing recall and precision for each mention m, even if m is a singleton. $B^{3}$ computes the intersection $|K_{i}\cap R_{j}|$ between every coreference chain $K_{i}$ in the key and every coreference chain $R_{j}$ in the response and then sums up recall and precision for each pair $i$, $j$ and normalizes. In turn, recall and precision for $i$, $j$ are computed by summing up recall and precision for each mention m in $|K_{i}\cap R_{j}|$. For instance, recall for $m$ is the proportion of mentions in $|K_{i}\cap R_{j}|$ and the number of mentions in $K_{i}$:
$\text{Recall}_{B^{3}}(m)=\frac{|K_{i}\cap R_{j}|}{|K_{i}|}\quad(m\in K_{i}\cap R_{j})$ (6)
Precision for $m$ is the proportion between $|K_{i}\cap R_{j}|$ and $|R_{j}|$.
3.3.3 An Entity-Based Metric: CEAF
$B^{3}$ also suffers from a problem—namely, that a single chain in the key or response can be credited several times. This leads to anomalies; for instance, if all coreference chains in the key are merged into one in the response, the $b^{3}$ recall is one. The ceaf metric was proposed by [24] to correct this problem. The key idea of CEAF is to align chains (entities) in the key and response using a map $g$ in such a way that each chain $K_{i}$ in the key is aligned with only one chain $g(K_{i})$ in the response and to then use the similarity $\phi(K_{i},g(K_{i}))$ to compute recall and precision. Because different maps are possible, the one that achieves optimal similarity is used. [24] present two entity-pair similarity measures that can be used in CEAF: one is the absolute number of common mentions between two entities, and the other is a “local” mention F-measure between two entities. The two measures lead to the mention-based CEAF-m and entity-based CEAF-e, respectively. [25]
4 Baseline Model & Extensions
We adopt a model from [1] and we use it as the baseline model (without higher-order inference). We use UDAPI [^3] https://github.com/udapi/udapi-python to load the data in CorefUD format. We use XLM-Roberta-large as an encoder.
Building on this baseline, we propose several extensions to the standard end-to-end model presented earlier in this paper. Our objective is to create a universal model suitable for all datasets in the CorefUD collection.
4.1 Cross-lingual Training
Some of the datasets are relatively small (see Table 1). To enhance performance, we propose pretraining the model on a concatenation of all training datasets within the CorefUD collection.
Table 4: Number of trainable parameters of the models
Model Pretrained params New params
mBERT 180M 40M
XLM-R 550M 50M
As you can see from Table 4, there are approximately 50 million parameters trained from scratch for XLM-R. For smaller datasets, it is practically impossible to train so many random parameters from scratch. To address this issue, we first pre-train the model on the joined dataset and then fine-tune the model for a specific language.
4.2 Syntactic Information
We believe that incorporating syntactic information can significantly enhance the model, particularly when manually annotated dependencies are available, such as in the Czech PDT dataset. Moreover, dependency information is essential for identifying mention heads, as some datasets in the CorefUD collection annotate mentions as nodes in the dependency tree rather than spans in the linear text flow.
To encode syntactic information, we add to each token representation its path to the ROOT in the dependency tree. In more detail, we first set the maximum tree depth parameter and then concatenate Bert representations of all parents up to max depth with the embedding of the corresponding dependency relation. Thus the resulting tree structure representation has the size of $max\_tree\_depth\times(bert\_emb\_size+deprel\_emb\_size)$. This representation is then concatenated with BERT embedding of each token.
4.3 Span2head Model
Given that the official scorer uses head-match evaluation, we decided to train the model to predict heads rather than entire spans to optimize the evaluation metric. Having all the useful information (even dependency trees), the model should learn the original rules for selecting the head.
A straightforward approach to get mention heads is to represent mention with its headword on the input. We describe this approach in detail later. However, this method is not ideal because multiple mentions can share the same head, which could result in merging distinct mentions and their clusters. To prevent this, we represent mentions with their full spans, predicting the head of each mention at the top of our model, and then outputting only the headword(s). This way, when building clusters, mentions are represented by their spans, ensuring that clusters of different mentions with the same head are not erroneously merged.
We implemented two versions of the head prediction model, both as separate classification heads on top of our coreference resolution model. The first model predicts the relative position of the headword(s) within a span using the hidden representation of the span from the coreference model. The output probabilities of head positions are obtained using sigmoid activation, allowing the model to predict multiple heads, even though only a single headword is present in the gold data. This serves as an optimization of the evaluation metric: if multiple words are likely headword candidates, it is statistically advantageous to output all of them.
Figure 3: First Span2Head Model
The second model uses a binary classification of each span and head candidate pair, so again, there can be more headwords of a single span predicted.
Figure 4: Second Span2Head Model
4.4 Head Representations
Later, inspired by the word-level coreference resolution model, we decided to try to model mentions only with their syntactic head. As mentioned above, For the CRAC official evaluation metric, reconstructing the original span is unnecessary, as predicting the correct head suffices. Therefore, we explored a simplified word-level coreference resolution approach that bypasses the span prediction step. Given that syntactic information is available in all CorefUD datasets, this is not so unrealistic scenario because in most cases the spans can be reconstructed by simply taking the whole subtree of the predicted node (head). Heads are heuristically selected only in cases where a mention does not form a single subtree. The proportion of such mentions is shown in Table 3(column non-tree). Using a word-level model reduces the mention space from quadratic to linear, making the model more efficient and minimizing potential false-positive mentions. Moreover, we believe that for very long mentions, the standard representation (sum of the start token, end token, and attended sum of all tokens) becomes insufficient.
4.5 Singletons
Some datasets in the CorefUD collection have singletons annotated and others do not. Specifically, in CorefUD 1.1, 8 out of 17 datasets have more than 10% singletons and 6 of these have more than 70% singletons which is probably a sign of consistent entity annotation independent of the coreference annotation. The baseline model completely ignores singletons during training [^4] The loss is the sum over all correct antecedents and since singletons have no gold antecedents they do not affect the loss.. As a result for these 6 singleton-including datasets, we discard more than 70% of training data for mention identification task. To leverage this data, we incorporate singleton modeling into our model and propose several approaches to address this.
4.5.1 Another Dummy Antecedent
In the first proposed approach, we introduce another virtual antecedent (with trained embedding) representing that the mention has no real antecedent, but it is a valid mention, so it is a singleton. For these mentions, we avoid using the binary score because it would unnecessarily model the similarity between all the singletons 555because all the singleton representations would be trained to be similar to the dummy antecedent. Additionally, included an option to use a separate feed-forward network for predicting singleton scores (different from the one for other mentions scores).
4.5.2 Mention Modeling
The second approach modifies the loss function to model mentions independently of coreference relations. In this approach, we simply add a binary cross-entropy of each span being a mention to the loss function. In other words, we add another classification head for the mention classification:
$J(D)=\log\prod_{i=1}^{N}\sum_{\hat{y}\in Y(i)\cap\texttt{GOLD}(i)}P(\hat{y})+\underbrace{y_{m}^{(i)}\cdot\sigma(s_{m}(i))+(1-y_{m}^{(i)})\cdot\sigma(-s_{m}(i))}_{\text{singletons binary cross-entropy}}$ (7)
where $y_{m}^{(i)}$ is $1$ if span $i$ corresponds to gold mention, 00 otherwise.
In the prediction step, the mention score is evaluated only for potential singletons. If a mention has no real antecedent we look at the mention score. If it is likely to be a mention we make it a singleton, otherwise it is not a mention at all.
4.6 Overlapping Segments
One limitation of the XLM-R model is the short sequence length. In the employed model, the input document is split into segments which are processed individually with a BERT-like encoder and merged in the antecedent-prediction head. To propagate errors correctly across segments, we need to process all segments in a single gradient update. However, due to GPU memory constraints, we set a maximum number of segments, and if a document exceeds this limit, we split it into multiple individual documents without any mutual coreferences.
We propose using segment overlapping with a cluster merging algorithm to address this issue. The cluster merging algorithm is straightforward: We iterate over all mentions in the new part and if it is present in a cluster in the previous part we simply make a union between its clusters from both parts. We experiment with minimal (just one overlapping segment) and maximal (each continuing example has just one new segment the rest is overlapping) segment overlap. For instance, if a document has 6 segments and the maximum number of segments for a single example is 4, then we split the document into 3 examples overlapping with 3 segments. The example is shown in Figure 5. We further extend this method by only using clusters in the continuing examples where at least one mention is present in new segments because if there is no such mention then all the mentions were seen already in the previous part with a longer left context.
Figure 5: Example of overlapping segments splitting
5 Experiments
5.1 Training
We trained all models on NVIDIA A40 GPUs using online learning (batch size 1 document). We limit the maximum sequence length to 6 non-overlapping segments of 512 tokens. During training, if the document is longer than $6\times 512$ tokens, a random segment offset is sampled to take a random continuous block of 6 segments, and the rest of them are discarded. During prediction, longer documents are split into independent sub-documents (for simplicity, non-overlapping again). For the head-only model, we double the maximum number of segments since the model is much more memory-effective therefore we can afford it. We train a model for each dataset for approximately 80k updates in our monolingual experiments. For joined-pre-trained models, we use 80k steps for model pre-training on all the datasets and approximately 30k for fine-tuning on each dataset. Each training took from 8 to 14 hours.
5.2 Baseline and Extensions
We begin by training a monolingual model for each language. For these experiments, we use XLM-Roberta-large and specific monolingual models for each language. Monolingual models for individual languages are listed in Table 5. Then we train a joined cross-lingual model for all the datasets which we fine-tune for each dataset afterwards.
To evaluate the effect of each extension, we first run all its variants described in 4 on the basic XLM-R-large model and measure the performance for each language (Sections 6.1.1 and 6.1.2). Then we compare each extension in its best configuration to the baseline joined model in Section 6.1.3 (since joined models achieve the best results from our three baseline models). In the last step, we evaluate the effects of Overlapping segments only on the best model configuration for each language (because it is orthogonal to all other extensions).
5.2.1 Monolingual Models
We evaluate a specific monolingual model for each language. We use a large variant where it is available. For Czech we use Czert [26], for English Roberta-large [27], gbert-large [28]. Other languages these models [29, 30, 31, 32, 33, 34, 35, 36]. The complete list can be found in Table 5.
5.3 Zero-shot Cross-lingual Experiments
We also conduct zero-shot cross-lingual experiments in two variants; dataset zero shot, and language zero shot. In dataset zero shot we train the model on all the datasets except one and then we evaluate the model on the excluded dataset. However, this approach is not a true cross-lingual zero-shot test, as multiple datasets exist for some languages, meaning the model was still exposed to the language of the evaluation data during training.
To address this, we performed language zero-shot experiments, where we removed all datasets for a particular language from the training data, and then evaluated the model on all datasets in that language. Additionally, we conducted a final experiment focusing on the ParCor corpora, which consists of parallel English and German datasets. This allowed us to perform a true zero-shot evaluation on both ParCor corpora.
6 Results & Discussion
Table 5 presents the results of baseline models on development data parts. The Monoling column shows the result of the monolingual model specific to each language. The XLM-R column presents results of XLM Roberta large trained separately for each dataset. The Joined column corresponds to the joined model described in Section 4.1. Several key observations can be drawn from these results. Most notably, monolingual models tend to outperform the cross-lingual XLM-Roberta model, especially for the Lithuanian and Czech PCEDT datasets, where the monolingual models—even though they are smaller—exhibit significantly better performance. The reason is probably in the difference of these two datasets from the other ones. We can see that for these two datasets, joined training does not help as much as for other datasets. For the rest of the datasets, joined training surpasses both other models (monolingual and multilingual trained from scratch).
Table 5: Results of baseline models on dev data. CRAC 2023 official evaluation metric
[…]
6.1 Effects of Extensions
6.1.1 Span2Head
Table 6 shows the performance of different variants of span2head model. None column shows the results of the baseline XLM-R-large model without any extensions. The Multi column reflects the results of the multi-class relative position classification (the first method described in Section 4.3). The binary column represents the second method (binary classification of span-headword candidate pair). We added Heads column with heads-only mention representation to show if it is better to use the span2head model or heads only model (since they are mutually exclusive). From the results, we can make several conclusions:
  1. 1.
    Span2head model slightly helps.
  2. 2.
    Binary model achieves better results, than the multi-class model.
  3. 3.
    For all the datasets where the span2head model helps, the heads-only model is much better so it does not make sense to use the span2head model in the next experiments.
Another interesting point is that for a monolingual model trained from scratch, the heads-only model improves the results on 13 out of 17 datasets and the results are worse only for a single dataset (English-ParCor).
Table 6: Results of XLM-R baseline with different types of Span2Head model
Dataset/S2H model None Multi Binary Heads
[…]
cs_pcedt 72.06 62.73 69.77 74.83
[…]
de_parcorfull 72.99 64.44 65.73 73.18
[…]
en_parcorfull 69.65 58.4 64.42 67.42
es_ancora 73.18 66.75 65.2 79.7
fr_democrat 55.97 50.18 51.3 60.78
hu_korkor 60.08 46.52 58.32 60.34
hu_szegedkoref 68.02 63.76 67.44 68.87
lt_lcc 72.48 69.5 72.01 75.13
no_bokmaalnarc 72.83 67.52 72.31 72.93
no_nynorsknarc 72.2 67.3 71.34 75.05
[…]
tr_itcc 20.51 17.32 17.08 27.82
avg 66.19 60.52 64.06 69.74
6.1.2 Singletons
Table 7 shows the results of all proposed variants of the singletons extension. Again, None column represents the baseline xlm-r-large trained from scratch for each dataset. Dummy column shows the results of plain dummy antecedent for singletons. mask extends the previous method binary score masking for singletons. This means that the method does not use the binary score for similarities to the singleton dummy antecedent. separate uses separate FFNN for singletons score prediction (different from the one used for standard mentions). The last method mentions represents the method described in section 4.5.2. Additionally, the Table shows the number of parameters added by each variant of the model. The last column shows the percentage of singleton entities in each dataset (datasets with few singletons do not need any singleton model).
Table 7: Results of XLM-R baseline with different types of singletons model
Model None Dummy Mask Separate Mentions % of singletons
+params 0 3092 3092 40M 40M -
ca_ancora 70.7 70.81 68.75 70.82 71.61 2.6
cs_pcedt 72.06 72.03 71.65 71.81 72.44 6.6
cs_pdt 71.21 70.75 70.04 72.23 72.84 40.8
de_parcorfull 72.99 70.23 70.09 65.47 71.29 6.2
de_potsdamcc 68.26 67.99 68.24 68.44 67.49 76.5
en_gum 64.06 70.4 69.81 70.89 71.61 73.8
en_parcorfull 69.65 67.49 64.78 65.89 64.43 6.9
es_ancora 73.18 72.88 70.98 73.74 73.85 3.3
fr_democrat 55.97 55.6 53.96 53.86 59.75 81.6
hu_korkor 60.08 60.56 56.6 55.59 57.91 0.9
hu_szegedkoref 68.02 67.1 66.57 67.1 67.03 8.0
lt_lcc 72.48 72.15 75.59 73.3 74.48 11.2
no_bokmaalnarc 72.83 73.28 73.33 73.84 74.15 89.4
no_nynorsknarc 72.2 72.99 73.57 74.77 74.26 88.7
pl_pcc 72.32 73.78 72.81 74.4 74.49 82.6
ru_rucor 68.68 67.56 67.91 68.55 69.12 3.3
tr_itcc 20.51 21.11 16.07 20.24 13.21 0.0
avg 66.19 66.38 65.34 65.94 66.47 -
From the table, we can see that on average the best results were achieved with the mention model. Mask model has very poor results, but it makes sense because the number of specific parameters for singletons modeling is very low (3 092 params is used as embedding for dummy antecedent) and if we discard the binary score, we reduce the learning power of the model even more. Dummy model does not help for any dataset statistically significantly. Separate and Mentions models are significantly better for Spanish (Mentions even for Catalan) which has almost no singletons. One reasonable explanation is that the baseline model is under-parametrized for this dataset. We evaluated this hypothesis by increasing the model size, but the improvement was not statistically significant. In the following chapter, we evaluate singletons on the model with joined pretraining, we compute confidence intervals, and in Table 13 we can see that the improvement is below statistical significance (on $95\%$ significance level). So we believe that these strange results in Table 7 are caused by random chance.
Another important conclusion from this table is that for all the datasets that have some singletons, the singletons model achieves better results than the baseline.
6.1.3 Best Combinations
Table 8 presents the performance of all the extensions proposed. The table also shows the 95% confidence intervals. Span2head and Only head extend Trees (both also uses syntax tree representations). We can see, that tree representations alone improve the results only for ParCor corpora. Span2Head model helps significantly only for English-GUM and Hungarian-korkor corpora but it is not better than only heads model. Singletons modeling helps for most of the datasets that have singletons annotated. The exceptions are in Norwegian Bokmaal and German Potsdam datasets. Another interesting point is that for German ParCor dataset singletons modeling helps significantly even though de-parcor itself does not contain singletons. One possible explanation is in the joined model. Adding the singletons from other languages to the training data can help to predict the mentions in German ParCor dataset. Results in Table 7 support this theory because we can see that when we train the model from scratch, singletons modeling does not help. Only heads model helps for almost all the datasets since it reduces the mention space and improves precision significantly. The extensions selected for the best model are highlighted in green color and the last column shows the results of the best combination. Please note that the results may show slight inconsistencies, even for datasets where a single extension was selected (e.g., ’Only heads’ and ’Best Comb.’ for Hungarian). This is because the experiments were rerun, averaging results across different runs with the same configuration. However, these differences are within the confidence intervals, ensuring the results remain reliable.
Table 8: Performance gains over the best base model with proposed extensions together with 95% confidence intervals. The models that surpass the baseline significantly are bold. The results which are better only on lower confidence level are underlined. Variants selected for the best models are green.
Dataset/Extension Joined Trees Span2head Singletons Only Heads Best Comb.
[…]
cs_pcedt 73.23 $\pm$ .08 73.2 $\pm$ .1 71.27 $\pm$ .11 73.17 $\pm$ .22 75.95 $\pm$ .13 75.89 $\pm$ .09
cs_pdt 73.47 $\pm$ .23 73.2 $\pm$ .46 72.98 $\pm$ .29 73.37 $\pm$ .28 79.43 $\pm$ .1 79.51 $\pm$ .02
[…]
fr_democrat 65.24 $\pm$ .49 65.46 $\pm$ .4 65.38 $\pm$ .49 66.23 $\pm$ .44 67.72 $\pm$ .81 68.58 $\pm$ .23
[…]
hu_szegedkoref 69.52 $\pm$ .59 69.59 $\pm$ .54 69.32 $\pm$ .34 69.47 $\pm$ .82 70.68 $\pm$ .46 70.67 $\pm$ .31
[…]
no_nynorsknarc 75.98 $\pm$ .17 74.79 $\pm$ .48 74.96 $\pm$ 1 76.59 $\pm$ .42 78.51 $\pm$ .21 78.81 $\pm$ .16
[…]
ru_rucor 70.38 $\pm$ .64 70.23 $\pm$ .42 70.49 $\pm$ .42 70.6 $\pm$ .41 75.8 $\pm$ .58 75.98 $\pm$ .42
tr_itcc 45.39 $\pm$ 1.3 34.79 $\pm$ 7.7 29.53 $\pm$ .5.2 45.63 $\pm$ .4 45.17 $\pm$ 3.9 44.17 $\pm$ 1.8
avg 70.95 70.94 69.69 71.77 73.70 74.67
6.1.4 Overlapping Segments
Table 9 evaluates longer context implementation through overlapping segments. We use a maximum length of 8 segments666since we use this maximum length in most of our other experiments, and we evaluate minimum and maximum overlap (1 and 7, respectively). We also try variants with filtering already-seen mentions as described in section 4.6 (columns min_filter and max_filter). For most of the datasets, the results do not change at all, so additionally, we run the same experiment with just 4 segments (Table 10)777which is the minimum number of segments we used in our models, and we count basic statistics of long coreference relations (Table 11). Table 11 presents three types of information:
  1. 1.
    column cross N segment corefs shows the percentage of coreference links that cross the boundary of N segments. These links cannot be predicted without the overlapping segments.
  2. 2.
    column nearest coref cross N is the percentage of mentions for which the nearest antecedent is already in a different segments block, so all the antecedents are in the different block and the antecedent cannot be predicted with the standard model.
  3. 3.
    column segments over N shows the percentage of segments that are after the first N-segments block. Only those segments beyond the first N-segment block are affected by predictions involving segment overlapping, as the remaining segments fit within the model’s capacity and are processed in a single example.
We can see that a significant number of log-distant coreferences are in French, Turkish, and Russian datasets. Some of them are also in both Czech datasets, Catalan and Polish. By far the most long coreferences are in the French dataset. From Table 9, we can see that segments overlapping decreases the performance. We investigated the dataset closely and found out that in this dataset multiple documents are merged into a single document quite often. When this happens, cross coreferences between those concatenated documents are not annotated so the model which can predict long-distant coreferences is penalized. To evaluate the models correctly, we add the variant where we split the documents in French dev data manually (row fr_split). After splitting, results with segments overlapping improve significantly, but they are still not as high as we expected in comparison with Russian, where the overlap helps significantly. To evaluate this a bit further, we also computed the recall for cross-segment coreference links. The recall is 72% for Russian and only 44% for French. By comparing these datasets manually, we conclude that the coreference chains are very long in the French dataset, and it is probably too hard to predict them correctly from the antecedents (one mistake in the chain might have a huge effect on the overall performance). For the Turkish dataset, it is hard to make any significant improvement due to its low quality.
Table 9: Results of the best model configurations with different approaches to merge segments during the prediction step. Maximum 8 segments
max. segments 8 None min min_filter max max_filter
[…]
cs_pcedt 78.23 $\pm$ .13 78.18 $\pm$ .11 78.23 $\pm$ .15 78.19 $\pm$ .15 78.12 $\pm$ .1
cs_pdt 80.02 $\pm$ .21 79.95 $\pm$ .19 79.97 $\pm$ .2 79.98 $\pm$ .2 79.96 $\pm$ .19
de_parcorfull 81.54 $\pm$ 1.3 81.54 $\pm$ 1.3 81.54 $\pm$ 1.3 81.54 $\pm$ 1.3 81.54 $\pm$ 1.3
de_potsdamcc 75.88 $\pm$ 2.2 75.88 $\pm$ 2.2 75.88 $\pm$ 2.2 75.88 $\pm$ 2.2 75.88 $\pm$ 2.2
en_gum 76.25 $\pm$ .12 76.25 $\pm$ .12 76.25 $\pm$ .12 76.25 $\pm$ .12 76.25 $\pm$ .12
en_parcorfull 67.92 $\pm$ 3.1 67.92 $\pm$ 3.1 67.92 $\pm$ 3.1 67.92 $\pm$ 3.1 67.92 $\pm$ 3.1
es_ancora 82.65 $\pm$ .27 82.65 $\pm$ .27 82.65 $\pm$ .27 82.65 $\pm$ .27 82.65 $\pm$ .27
[…]
hu_korkor 73.41 $\pm$ .43 73.41 $\pm$ .43 73.41 $\pm$ .43 73.41 $\pm$ .43 73.41 $\pm$ .43
hu_szegedkoref 71.2 $\pm$ .14 71.2 $\pm$ .14 71.2 $\pm$ .14 71.2 $\pm$ .14 71.2 $\pm$ .14
lt_lcc-corefud 76.82 $\pm$ .41 76.82 $\pm$ .41 76.82 $\pm$ .41 76.82 $\pm$ .41 76.82 $\pm$ .41
no_bokmaalnarc 78.96 $\pm$ .37 78.96 $\pm$ .37 78.96 $\pm$ .37 78.96 $\pm$ .37 78.96 $\pm$ .37
no_nynorsknarc 80.5 $\pm$ .14 80.5 $\pm$ .14 80.5 $\pm$ .14 80.5 $\pm$ .14 80.5 $\pm$ .14
[…]
ru_rucor 75.27 $\pm$ .42 77.23 $\pm$ .63 76.65 $\pm$ .39 76.77 $\pm$ .39 77.19 $\pm$ .57
[…]
avg 74.46 74.60 74.60 74.57 74.60
Table 10: Results of the best model configurations with different approaches to merge segments during the prediction step. Maximum 4 segments
max. segments 4 none min min filter max max filter
[…]
cs_pcedt 77.82 $\pm$ .088 77.29 $\pm$ .17 77.47 $\pm$ .17 77.84 $\pm$ .096 77.65 $\pm$ .11
cs_pdt 79.9 $\pm$ .22 79.87 $\pm$ .16 79.84 $\pm$ .2 79.91 $\pm$ .22 79.74 $\pm$ .19
de_parcorfull 81.54 $\pm$ 1.3 81.54 $\pm$ 1.3 81.54 $\pm$ 1.3 81.54 $\pm$ 1.3 81.54 $\pm$ 1.3
de_potsdamcc 75.88 $\pm$ 2.2 75.88 $\pm$ 2.2 75.88 $\pm$ 2.2 75.88 $\pm$ 2.2 75.88 $\pm$ 2.2
en_gum 76.25 $\pm$ .12 76.25 $\pm$ .12 76.25 $\pm$ .12 76.25 $\pm$ .12 76.25 $\pm$ .12
en_parcorfull 67.92 $\pm$ 3.1 67.92 $\pm$ 3.1 67.92 $\pm$ 3.1 67.92 $\pm$ 3.1 67.92 $\pm$ 3.1
es_ancora 82.59 $\pm$ .27 82.48 $\pm$ .31 82.56 $\pm$ .26 82.59 $\pm$ .27 82.57 $\pm$ .28
fr_democrat 68.81 $\pm$ .84 68.64 $\pm$ .89 68.33 $\pm$ .79 68.81 $\pm$ .92 67.77 $\pm$ 1.7
[…]
hu_korkor 73.41 $\pm$ .43 73.41 $\pm$ .43 73.41 $\pm$ .43 73.41 $\pm$ .43 73.41 $\pm$ .43
hu_szegedkoref 71.2 $\pm$ .14 71.2 $\pm$ .14 71.2 $\pm$ .14 71.2 $\pm$ .14 71.2 $\pm$ .14
lt_lcc 76.82 $\pm$ .41 76.82 $\pm$ .41 76.82 $\pm$ .41 76.82 $\pm$ .41 76.82 $\pm$ .41
no_bokmaalnarc 78.96 $\pm$ .36 78.91 $\pm$ .38 78.87 $\pm$ .34 78.96 $\pm$ .36 78.47 $\pm$ .37
no_nynorsknarc 79.87 $\pm$ .57 80.32 $\pm$ .18 80.15 $\pm$ .37 79.87 $\pm$ .57 78.94 $\pm$ .12
pl_pcc 75.94 $\pm$ .15 76.05 $\pm$ .18 76.03 $\pm$ .22 75.93 $\pm$ .15 75.71 $\pm$ .17
ru_rucor 75.58 $\pm$ .37 76.54 $\pm$ .31 75.89 $\pm$ .55 75.62 $\pm$ .42 72.8 $\pm$ .25
tr_itcc 43.61 $\pm$ 1.5 43.85 $\pm$ 1.8 43.81 $\pm$ 1.3 43.59 $\pm$ 1.5 42.19 $\pm$ 2.7
avg 74.31 74.39 74.33 74.24 73.88
Table 11: Statistics of long documents and distant coreference relations across the datasets
cross N segment corefs
nearest coref cross N
segments over N
4 8 4 8 4 8
ca_ancora 5.975 1.459 .6044 .2747 4.739 1.422
cs_pcedt 15.82 7.398 1.107 .1441 13.55 2.133
cs_pdt 9.787 1.54 .4787 .06769 4.175 .3976
de_parcorfull 0.0 0.0 0.0 0.0 0.0 0.0
de_potsdamcc 0.0 0.0 0.0 0.0 0.0 0.0
en_gum 0.0 0.0 0.0 0.0 0.0 0.0
en_parcorfull 0.0 0.0 0.0 0.0 0.0 0.0
es_ancora 3.012 0.0 .1396 0.0 .5076 0.0
fr_democrat 52.86 25.12 1.664 .8106 34.86/27.52 27.52/12.84
hu_korkor 0.0 0.0 0.0 0.0 0.0 0.0
hu_szegedkoref 0.0 0.0 0.0 0.0 0.0 0.0
lt_lcc 0.0 0.0 0.0 0.0 0.0 0.0
no_bokmaalnarc 9.026 0.0 .1092 0.0 3.846 0.0
no_nynorsknarc 15.49 0.0 .5405 0.0 10.29 0.0
pl_pcc 11.79 2.724 .3591 .1554 2.954 .8439
ru_rucor 43.77 21.76 3.273 1.283 35.44 16.46
tr_itcc 55.6 25.07 5.703 2.037 55.56 11.11
6.2 Cross-lingual Transfer
Table 12: Zero-shot cross-lingual transfer results. The table compares four settings. A joined model trained for all datasets, a joined model fine-tuned on each dataset separately afterwards, zero-shot setting for each dataset, and zero-shot setting for each language.
Dataset Zero-shot Finetunned Joined Zero-shot Lang. Zero-shot
ca_ancora 82.12 $\pm$ .13 81.9225 $\pm$ 0.38 72.01 $\pm$ .17 58.53 $\pm$ .49
cs_pcedt 75.89 $\pm$ .09 74.7775 $\pm$ 0.23 62.26 $\pm$ .02 51.78 $\pm$ .24
cs_pdt 79.51 $\pm$ .02 78.8575 $\pm$ 0.25 69.92 $\pm$ .16 64.55 $\pm$ .28
[…]
de_potsdamcc 78.41 $\pm$ .89 75.1325 $\pm$ 2.24 67.15 $\pm$ 1.4 64.33 $\pm$ .64
[…]
es_ancora 82.43 $\pm$ .07 82.245 $\pm$ 0.35 75.09 $\pm$ .26 61.68 $\pm$ .38
fr_democrat 68.58 $\pm$ .23 67.7225 $\pm$ 0.5 60.32 $\pm$ .14 -
hu_korkor 73.55 $\pm$ .57 73.2625 $\pm$ 0.54 58.26 $\pm$ .31 53.75 $\pm$ .35
hu_szegedkoref 70.67 $\pm$ .31 69.75 $\pm$ 0.65 55.16 $\pm$ .11 53.36 $\pm$ .17
lt_lcc 77.65 $\pm$ .71 75.65 $\pm$ 0.78 47.93 $\pm$ .59 -
no_bokmaalnarc 78.15 $\pm$ .29 77.8225 $\pm$ 0.67 74.83 $\pm$ .05 64.77 $\pm$ .76
no_nynorsknarc 78.81 $\pm$ .16 77.955 $\pm$ 0.46 72.7 $\pm$ .37 63.24 $\pm$ .94
pl_pcc 76 $\pm$ .17 76.0625 $\pm$ 0.08 58.44 $\pm$ .13 -
ru_rucor 75.98 $\pm$ .42 72.665 $\pm$ 0.83 62.12 $\pm$ .28 -
tr_itcc 44.17 $\pm$ 1.8 38.52 $\pm$ 1.61 46.26 $\pm$ .2 -
avg 74.67 71.09 62.19 58.92
Table 12 shows the results of cross-lingual transfer evaluation. We can see that for most of the datasets fine-tuning the model on the specific data helps but the difference is not so significant. For several datasets, the results are even comparable. For small datasets, the fine-tuned models achieve much better results than the joined model. Note that we do not use any weighing of particular datasets in the joined collection so small datasets have a marginal effect on the training. The results of dataset zero-shot and language zero-shot are mostly very predictable. For the languages that have more datasets in CorefUD collection, the drop of zero-shot compared to the joined model was smaller than for the rest of the datasets. When we discard all the datasets for each language, the results are mostly significantly worse compared to discarding just single dataset. This behavior is expected. Several exceptions should be analyzed more. The strangest results are on the Turkish dataset. Here the zero-shot model achieves the best performance. Together with the fact that the numbers are very low for this dataset, it suggests that there is some noise in the dataset annotation. If we train the model on the training part of the Turkish dataset the noise is present on two sides (training and testing). If we train a zero-shot model we eliminate one part.
Another interesting aspect is the superior performance of de-parcor and english-gum dataset in the language zero-shot scenario
6.3 Final Results
Table 13 presents the final results of the best-performing models on a test set. For comparison, we use the best model of CRAC 2023 Multilingual coreference resolution shared task from [37] but not the best submission. The best model is much larger than ours, and it uses ensembling, so as the main model for comparison, we use their RemBERT version without ensembling, which has a comparable size to our model.
From the table, we can see that our model outperformed CorPipe on most of the datasets. For some datasets (mostly the ones without singletons annotated), it outperformed even the large CorPipe model, which has approximately 3 times more trainable parameters. The only dataset where CorPipe outperformed our model by a large margin is the Polish one, which we consider an anomaly because, for this dataset, CorPipe results on the test set are much better than for the dev set. On the dev set, we achieve similar results. The same happened in the opposite way for the Russian dataset. Surprisingly, we even outperformed the large CorPipe model in average score, but this is caused mainly by a large margin on Turkish and German-parcor.
Table 13: Final results on test sets
Our Best CorPipe CorPipe-large
ca_ancora 82.05 79.93 82.39
[…]
de_potsdamcc 71.07 72.63 67.93
en_gum 73.19 72.33 75.02
en_parcorfull 61.62 57.58 64.79
[…]
hu_szegedkoref 69.02 65.4 69.16
lt_lcc 70.69 68.63 75.63
no_bokmaalnarc 76.24 75.43 78.94
no_nynorsknarc 74.54 73.64 77.24
[…]
7 Conclusion
In this paper, we explored and evaluated various approaches to multilingual coreference resolution using the CorefUD 1.1 dataset. Our experiments revealed that monolingual models typically outperform cross-lingual models, especially for languages with datasets that are distinct in their characteristics. However, joint training across languages can still provide benefits for most of the datasets where sufficient cross-linguistic similarities exist.
We proposed several extensions to enhance the baseline models, including cross-lingual training, Span2Head modeling, syntactic information integration, headword mention representation, and long-context prediction. Among these, the heads-only model and singleton modeling showed the most consistent improvements across different datasets, demonstrating the importance of targeted adaptations for coreference resolution tasks. For several datasets, long-context prediction brings also a significant improvement. Additionally, our zero-shot cross-lingual experiments provided insights into the challenges and opportunities of cross-lingual transfer, with the Turkish dataset results highlighting potential issues related to noise in the training data.
Overall, our findings emphasize the need for tailored approaches in coreference resolution, particularly when dealing with diverse languages and annotation schemes. Future work could further investigate the impact of dataset-specific characteristics on model performance and explore additional strategies for enhancing cross-lingual transfer, especially for low-resource languages. Our source codes are publicly available for subsequent research [^8] https://github.com/ondfa/coref-multiling.
Acknowledgement
Computational resources were supplied by the project “e-Infrastruktura CZ” (e-INFRA LM2018140) provided within the program Projects of Large Research, Development and Innovations Infrastructures. This work has been supported by Grant No. SGS-2022-016 Advanced methods of data processing and analysis.